# Ordinal Adaptive Correction: A Data-Centric Approach to Ordinal Image Classification with Noisy Labels

### v1

@@ Abstract
I Introduction
@@
The performance, fairness, and robustness of modern artificial intelligence (AI) systems are fundamentally shaped by the quality of the data they are trained on. This has given rise to a data-centric view of AI, which posits that systematic improvements in data quality are often more impactful than architectural innovations alone [1]. In computer vision, the success of deep neural networks (DNNs) has been fueled by large-scale annotated datasets like ImageNet [2]. However, creating such datasets is a costly and labor-intensive process. Consequently, researchers often rely on scalable but less controlled methods like crowdsourcing or web scraping, which inevitably introduce noisy or incorrect labels into the training data [3].
@@
The prevalence of noisy labels in benchmark datasets such as ImageNet [2], CIFAR-10 [4], and MNIST [5] presents a significant challenge. DNNs trained on such data can exhibit severely degraded performance, as they may memorize the incorrect annotations. This raises a fundamental question: how can we effectively train models in the presence of label noise? This problem is particularly acute in ordinal classification (or ordinal regression), where the goal is to predict a label from a set of classes with an inherent order, such as estimating age, grading disease severity, or ranking customer satisfaction. In these tasks, the semantic closeness of adjacent classes makes annotation inherently ambiguous and increases the likelihood of label noise.
@@
Existing approaches to mitigate label noise fall into two main categories. Model-centric methods aim to make the learning process itself robust, for example, by designing noise-tolerant loss functions or using regularization to prevent overfitting to incorrect labels [6]. In contrast, data-centric approaches focus on the data itself, with the dominant strategy being sample selection. Methods like CASSOR [7] and ICDF [8] identify and filter out samples that are likely mislabeled, training the model on a cleaner subset. While effective, the primary drawback of this approach is the outright removal of samples, which discards the potentially valuable information contained within the features of those instances. This reveals a significant research gap: a need for methods that can correct rather than discard noisy labels by adaptively modeling the uncertainty inherent in ordinal data.
@@
To address this gap, this paper introduces a novel framework named ORDinal Adaptive Correction (ORDAC). Instead of discarding samples suspected of having noisy labels, ORDAC is designed to correct them by leveraging the expressive power of Label Distribution Learning (LDL) [9]. The core idea is to represent each label not as a single value but as a Gaussian distribution, where the mean represents the label value and the standard deviation quantifies the uncertainty. The ORDAC framework operates iteratively, using the model’s own predictions in a cross-validation setup to dynamically update the mean and standard deviation of each training sample’s label distribution. This adaptive correction mechanism allows the model to gradually learn from progressively cleaner and more reliable labels, enhancing its robustness and generalization.
@@ II Related Work
II-A Learning from Noisy Labels
@@
The paradigm of Data-Centric AI emphasizes that improving data quality is a cornerstone of building robust models [10]. A central challenge within this paradigm is learning from data with noisy labels. The strategies developed to address this can be broadly grouped into two families.
@@
Model-centric approaches aim to make the learning algorithm itself resilient to noise. This includes designing robust loss functions like Generalized Cross-Entropy [11] that are less sensitive to large errors, or employing regularization techniques like Mixup [12] that discourage the model from memorizing incorrect labels. While often effective, these methods treat the dataset as a static, unchangeable entity.
@@
Data-centric approaches, in contrast, focus on improving the dataset itself [3]. These methods generally fall into two categories: sample selection and label correction. The more common strategy has been sample selection, which aims to detect and reduce the effect of mislabeled samples. Many early works relied on the small-loss trick, based on the observation that deep networks tend to learn clean samples earlier. This idea underpins methods such as Co-teaching [13], where two networks exchange small-loss samples, and MentorNet [14], which learns a curriculum to prioritize easier, likely clean samples. Over time, this paradigm has evolved into label quality scoring frameworks. For example, Confident Learning [15] provides a principled, model-agnostic way to identify label errors by analyzing the relationship between noisy and predicted labels. Similarly, model-agnostic label quality scoring [16] introduces multiple scoring metrics to detect low-quality samples, and Dataset Cartography [17] visualizes learning dynamics to identify hard-to-learn samples that often correspond to mislabeled data. These approaches have proven effective not only in image classification but also in more complex tasks such as object detection [18] and semantic segmentation [19].
@@
In contrast, label correction offers a more direct, data-restorative strategy. Rather than discarding noisy samples, these methods try to infer the correct labels, preserving the full dataset for training. Although promising, this line of work remains less explored compared to sample selection. Some methods adopt a meta-learning framework, where a meta-model learns a correction function for the noisy dataset, as in Meta Label Correction [20]. Others exploit the geometric structure of the feature space, for example, by using graph-based label propagation [21] to refine labels. Despite their strengths, most correction methods are designed for nominal classification tasks. Our work, ORDAC, advances this area by introducing a new label correction mechanism specifically tailored to the unique structural properties of ordinal data.
II-B Ordinal Classification and Its Challenges
@@
Ordinal classification, or ordinal regression, addresses supervised learning tasks where labels possess a natural order, such as age estimation or clinical severity grading [22]. Ignoring this intrinsic order and treating the problem as nominal classification is suboptimal, as the cost of misclassification is not uniform; for instance, a one-rank error is far less severe than a five-rank error.
@@
To leverage this structure, a diverse set of methodologies has been developed. A prominent family of methods is based on ordinal binary decomposition, which reframes the $K$-rank problem into a series of simpler binary classification tasks (e.g., is the rank >k>k?). Early deep learning approaches like OR-CNN [23] applied this principle, but a key challenge was ensuring that the binary predictions were monotonically consistent. The CORAL framework [24] elegantly solved this by sharing weights across the binary classifiers, a technique further refined by its successor, CORN [25]. Other major paradigms include threshold models, which learn a mapping to a latent continuous score that is then partitioned by a set of learned thresholds [26, 27], and the direct design of ordinal loss functions that explicitly penalize predictions based on their rank distance from the true label, such as the Weighted Kappa Loss [28]. While powerful, these methods are typically designed for clean datasets and their performance can degrade significantly in the presence of label noise, as they lack an explicit mechanism to handle incorrect rank annotations.
II-C Label Distribution Learning for Ordinal Tasks
@@
Label Distribution Learning (LDL) has emerged as a particularly well-suited paradigm for ordinal tasks due to its ability to handle label ambiguity [9]. Instead of a single hard label, LDL assigns a probability distribution over all possible labels to each instance. For ordinal data, this is often a unimodal distribution (e.g., a Gaussian), which naturally captures the semantic closeness of adjacent classes and the decreasing likelihood of distant ranks.
@@
LDL methods can be categorized as Fixed-form (FLDL) or Adaptive (ALDL) [29]. FLDL methods like DLDL-v2 [30] assume a static shape for the label distribution (e.g., a Gaussian with a fixed standard deviation). This is a strong assumption that limits the model’s capacity to represent varying levels of uncertainty across different samples. ALDL methods, such as those using a Unimodal-Concentrated Loss [29] or modeling ordinal relationships explicitly [31], offer more flexibility. However, it is crucial to note that the reported gains of many specialized loss functions have been questioned by studies highlighting the disproportionate impact of evaluation protocols, such as inconsistent data splitting, on final performance [32]. While our framework initializes with a fixed-form distribution, characteristic of FLDL, its core novelty lies in a dynamic correction mechanism that adaptively modifies each sample’s distribution, thus operating as an ALDL method.
II-D Learning Ordinal Regression with Noisy Labels
@@
The specific challenge of learning ordinal regression from noisy labels is an emerging research frontier. The few existing methods are primarily data-centric and have focused on sample selection. CASSOR [7], for example, proposes a class-aware selection strategy that estimates an ”insufficiency score” to dynamically adjust the sampling rate for each class. Similarly, ICDF [8] introduces a filtering algorithm based on inter-class feature-space distances to identify and remove noisy samples.
@@ III-A Framework Overview
    Iterative Correction Phase: After the warm-up phase, for each subsequent epoch, the adaptive correction mechanism is activated. The label distribution for each sample in the training set is updated based on the predictions from the corresponding model that held it out as a validation sample. For instance, the corrected labels from $D_{k}^{\text{valid}}$ (generated by model $M_{k}$) are used to update the corresponding samples in all other training sets $D_{j}^{\text{train}}$ where $j\neq k$. This process is detailed in Section III-B and Algorithm 1.
III-B Adaptive Label Correction Mechanism
@@
III-B1 Class-wise Prediction Debiasing
@@
III-B2 Sample-wise Distribution Update
@@ IV Results
TABLE I: Main results comparing our proposed methods (ORDAC and its variants) against baselines on the Adience and DR datasets under different noise rates ($\tau$). We report Macro-Averaged MAE and Recall (REC). Lower MAE and higher REC are better. Best results are in bold, second best are underlined.
Dataset $\tau$ Metric CORAL [24] DLDL-v2 [30] ORDAC ORDACC ORDACR
@@
IV-A Experimental Setup
IV-A1 Datasets
@@
  • Adience [33]: This dataset contains approximately 26K unfiltered face images from Flickr albums, annotated for age and gender. We focus on the age estimation task, which consists of 8 ordinal classes: (0-2), (4-6), (8-13), (15-20), (25-32), (38-43), (48-53), and (60+). The dataset’s in-the-wild nature, with variations in lighting, pose, and resolution, makes it a challenging benchmark. Crucially, the labels are derived from user profiles and are known to contain inherent noise due to self-reporting errors or mismatches between the uploader and the person in the photo. We use the aligned version of the dataset and the official 5-fold split as recommended by [32].
  • Diabetic Retinopathy (DR) [34]: This medical imaging dataset contains over 88K high-resolution retinal fundus images for detecting and grading diabetic retinopathy. The task is to classify each image into one of 5 ordinal severity levels: (0: No DR, 1: Mild, 2: Moderate, 3: Severe, 4: Proliferative DR). The dataset exhibits significant challenges, including variations in imaging conditions and image quality. We use the official training/test split and further divide the training data into 5 folds for our cross-validation setup.
IV-A2 Evaluation Metrics
@@
IV-A3 Noise Injection Protocol
Label noise in ordinal tasks is typically asymmetric, as mislabelings are more likely to occur between adjacent, semantically similar classes. To simulate this realistically, we inject Gaussian Asymmetric Noise into the training and validation sets, following the protocol used in prior works [35, 7, 8]. A noise transition matrix $T$, where $T_{ij}=P(\tilde{y}=j|y=i)$ is the probability of a true label $i$ being flipped to a noisy label $j$, is generated. The off-diagonal elements are defined by a Gaussian function of the distance between ranks:
@@ IV-A3 Noise Injection Protocol
IV-A4 Implementation Details
All experiments use a ResNet-50 architecture, pre-trained on ImageNet, as the feature extractor backbone. For our proposed method, the model is trained for a maximum of 50 epochs ($E_{\text{max}}$), with the adaptive correction mechanism activating after 10 warm-up periods ($E_{\text{corr}}$). The base learning rates for correction were set to $\alpha_{\text{base}}=0.2$ and $\beta_{\text{base}}=0.8$ based on the hyperparameter analysis in IV-E2.
IV-B Performance under Label Noise
We first evaluate the robustness of ORDAC against synthetic label noise. We compare our three proposed variants (ORDAC, ORDACC, and ORDACR) with two strong baselines: CORAL [24], a standard ordinal regression method, and DLDL-v2 [30], a fixed-form LDL method. The training and validation sets are corrupted with Gaussian asymmetric noise at rates $\tau\in\{0.2,0.4\}$. The case $\tau=0$ corresponds to training on the original, uncorrupted datasets.
The results are presented in Table I and visualized in Figure 2. Our proposed methods consistently and significantly outperform the baselines across both datasets and all noise levels. As expected, the performance of all methods degrades as the noise rate increases. However, the ORDAC framework demonstrates substantially greater resilience. For instance, on Adience with 40% noise, ORDACR achieves an MAE of 0.6283, a dramatic improvement over CORAL (1.0427) and DLDL-v2 (0.8649).

### v2

@@ Abstract
journal: Neurocomputing
[…]
[inst1]organization=School of Computer Engineering, Iran University of Science and Technology, country=Islamic Republic of Iran
1 Introduction
@@
The performance, fairness, and robustness of modern artificial intelligence (AI) systems are fundamentally shaped by the quality of the data they are trained on. This has given rise to a data-centric view of AI, which posits that systematic improvements in data quality are often more impactful than architectural innovations alone [sambasivan2021everyone]. In computer vision, the success of deep neural networks (DNNs) has been fueled by large-scale annotated datasets like ImageNet [russakovsky2015imagenet]. However, creating such datasets is a costly and labor-intensive process. Consequently, researchers often rely on scalable but less controlled methods like crowdsourcing or web scraping, which inevitably introduce noisy or incorrect labels into the training data [cordeiro2020survey].
@@
The prevalence of noisy labels in benchmark datasets such as ImageNet [russakovsky2015imagenet], CIFAR-10 [krizhevsky2009learning], and MNIST [deng2012mnist] presents a significant challenge. DNNs trained on such data can exhibit severely degraded performance, as they may memorize the incorrect annotations. This raises a fundamental question: how can we effectively train models in the presence of label noise? This problem is particularly acute in ordinal classification (or ordinal regression), where the goal is to predict a label from a set of classes with an inherent order, such as estimating age, grading disease severity, or ranking customer satisfaction. In these tasks, the semantic closeness of adjacent classes makes annotation inherently ambiguous and increases the likelihood of label noise.
@@
Existing approaches to mitigate label noise fall into two main categories. Model-centric methods aim to make the learning process itself robust, for example, by designing noise-tolerant loss functions or using regularization to prevent overfitting to incorrect labels [song2022learning]. In contrast, data-centric approaches focus on the data itself, with the dominant strategy being sample selection. Methods like CASSOR [yuan2023cassor] and ICDF [jiang2024noise] identify and filter out samples that are likely mislabeled, training the model on a cleaner subset. While effective, the primary drawback of this approach is the outright removal of samples, which discards the potentially valuable information contained within the features of those instances. This reveals a significant research gap: a need for methods that can correct rather than discard noisy labels by adaptively modeling the uncertainty inherent in ordinal data.
@@
To address this gap, this paper introduces a novel framework named ORDinal Adaptive Correction (ORDAC). Instead of discarding samples suspected of having noisy labels, ORDAC is designed to correct them by leveraging the expressive power of Label Distribution Learning (LDL) [geng2016label]. The core idea is to represent each label not as a single value but as a Gaussian distribution, where the mean represents the label value and the standard deviation quantifies the uncertainty. The ORDAC framework operates iteratively, using the model’s own predictions in a cross-validation setup to dynamically update the mean and standard deviation of each training sample’s label distribution. This adaptive correction mechanism allows the model to gradually learn from progressively cleaner and more reliable labels, enhancing its robustness and generalization.
@@ 2 Related Work
2.1 Learning from Noisy Labels
@@
The paradigm of Data-Centric AI emphasizes that improving data quality is a cornerstone of building robust models [zha2023data]. A central challenge within this paradigm is learning from data with noisy labels. The strategies developed to address this can be broadly grouped into two families.
@@
Model-centric approaches aim to make the learning algorithm itself resilient to noise. This includes designing robust loss functions like Generalized Cross-Entropy [zhang2018generalized] that are less sensitive to large errors, or employing regularization techniques like Mixup [zhang2017mixup] that discourage the model from memorizing incorrect labels. While often effective, these methods treat the dataset as a static, unchangeable entity.
@@
Data-centric approaches, in contrast, focus on improving the dataset itself [cordeiro2020survey]. These methods generally fall into two categories: sample selection and label correction. The more common strategy has been sample selection, which aims to detect and reduce the effect of mislabeled samples. Many early works relied on the small-loss trick, based on the observation that deep networks tend to learn clean samples earlier. This idea underpins methods such as Co-teaching [han2018co], where two networks exchange small-loss samples, and MentorNet [jiang2018mentornet], which learns a curriculum to prioritize easier, likely clean samples. Over time, this paradigm has evolved into label quality scoring frameworks. For example, Confident Learning [northcutt2021confident] provides a principled, model-agnostic way to identify label errors by analyzing the relationship between noisy and predicted labels. Similarly, model-agnostic label quality scoring [kuan2022model] introduces multiple scoring metrics to detect low-quality samples, and Dataset Cartography [swayamdipta2020dataset] visualizes learning dynamics to identify hard-to-learn samples that often correspond to mislabeled data. These approaches have proven effective not only in image classification but also in more complex tasks such as object detection [tkachenko2023objectlab] and semantic segmentation [lad2023estimating].
@@
In contrast, label correction offers a more direct, data-restorative strategy. Rather than discarding noisy samples, these methods try to infer the correct labels, preserving the full dataset for training. Although promising, this line of work remains less explored compared to sample selection. Some methods adopt a meta-learning framework, where a meta-model learns a correction function for the noisy dataset, as in Meta Label Correction [zheng2021meta]. Others exploit the geometric structure of the feature space, for example, by using graph-based label propagation [zhang2021dualgraph] to refine labels. Despite their strengths, most correction methods are designed for nominal classification tasks. Our work, ORDAC, advances this area by introducing a new label correction mechanism specifically tailored to the unique structural properties of ordinal data.
2.2 Ordinal Classification and Its Challenges
@@
Ordinal classification, or ordinal regression, addresses supervised learning tasks where labels possess a natural order, such as age estimation or clinical severity grading [gutierrez2015ordinal]. Ignoring this intrinsic order and treating the problem as nominal classification is suboptimal, as the cost of misclassification is not uniform; for instance, a one-rank error is far less severe than a five-rank error.
@@
To leverage this structure, a diverse set of methodologies has been developed. A prominent family of methods is based on ordinal binary decomposition, which reframes the $K$-rank problem into a series of simpler binary classification tasks (e.g., is the rank >k>k?). Early deep learning approaches like OR-CNN [niu2016ordinal] applied this principle, but a key challenge was ensuring that the binary predictions were monotonically consistent. The CORAL framework [cao2020rank] elegantly solved this by sharing weights across the binary classifiers, a technique further refined by its successor, CORN [shi2023deep]. Other major paradigms include threshold models, which learn a mapping to a latent continuous score that is then partitioned by a set of learned thresholds [vargas2020cumulative, herbrich1999support], and the direct design of ordinal loss functions that explicitly penalize predictions based on their rank distance from the true label, such as the Weighted Kappa Loss [de2018weighted]. While powerful, these methods are typically designed for clean datasets and their performance can degrade significantly in the presence of label noise, as they lack an explicit mechanism to handle incorrect rank annotations.
2.3 Label Distribution Learning for Ordinal Tasks
@@
Label Distribution Learning (LDL) has emerged as a particularly well-suited paradigm for ordinal tasks due to its ability to handle label ambiguity [geng2016label]. Instead of a single hard label, LDL assigns a probability distribution over all possible labels to each instance. For ordinal data, this is often a unimodal distribution (e.g., a Gaussian), which naturally captures the semantic closeness of adjacent classes and the decreasing likelihood of distant ranks.
@@
LDL methods can be categorized as Fixed-form (FLDL) or Adaptive (ALDL) [li2022unimodal]. FLDL methods like DLDL-v2 [gao2018age] assume a static shape for the label distribution (e.g., a Gaussian with a fixed standard deviation). This is a strong assumption that limits the model’s capacity to represent varying levels of uncertainty across different samples. ALDL methods, such as those using a Unimodal-Concentrated Loss [li2022unimodal] or modeling ordinal relationships explicitly [wen2023ordinal], offer more flexibility. However, it is crucial to note that the reported gains of many specialized loss functions have been questioned by studies highlighting the disproportionate impact of evaluation protocols, such as inconsistent data splitting, on final performance [paplham2024call]. While our framework initializes with a fixed-form distribution, characteristic of FLDL, its core novelty lies in a dynamic correction mechanism that adaptively modifies each sample’s distribution, thus operating as an ALDL method.
2.4 Learning Ordinal Regression with Noisy Labels
@@
The specific challenge of learning ordinal regression from noisy labels is an emerging research frontier. The few existing methods are primarily data-centric and have focused on sample selection. CASSOR [yuan2023cassor], for example, proposes a class-aware selection strategy that estimates an "insufficiency score" to dynamically adjust the sampling rate for each class. Similarly, ICDF [jiang2024noise] introduces a filtering algorithm based on inter-class feature-space distances to identify and remove noisy samples.
@@ 3.1 Framework Overview
    Iterative Correction Phase: After the warm-up phase, for each subsequent epoch, the adaptive correction mechanism is activated. The label distribution for each sample in the training set is updated based on the predictions from the corresponding model that held it out as a validation sample. For instance, the corrected labels from $D_{k}^{\text{valid}}$ (generated by model $M_{k}$) are used to update the corresponding samples in all other training sets $D_{j}^{\text{train}}$ where $j\neq k$. This process is detailed in Section 3.2 and Algorithm 1.
3.2 Adaptive Label Correction Mechanism
@@
3.2.1 Class-wise Prediction Debiasing
@@
3.2.2 Sample-wise Distribution Update
@@ 4 Results
Table 1: Main results comparing our proposed methods (ORDAC and its variants) against baselines on the Adience and DR datasets under different noise rates ($\tau$). We report Macro-Averaged MAE and Recall (REC). Lower MAE and higher REC are better. Best results are in bold, second best are underlined.
Dataset $\tau$ Metric CORAL [cao2020rank] DLDL-v2 [gao2018age] ORDAC ORDACC ORDACR
@@
4.1 Experimental Setup
4.1.1 Datasets
@@
  • 1.
    Adience [eidinger2014age]: This dataset contains approximately 26K unfiltered face images from Flickr albums, annotated for age and gender. We focus on the age estimation task, which consists of 8 ordinal classes: (0-2), (4-6), (8-13), (15-20), (25-32), (38-43), (48-53), and (60+). The dataset’s in-the-wild nature, with variations in lighting, pose, and resolution, makes it a challenging benchmark. Crucially, the labels are derived from user profiles and are known to contain inherent noise due to self-reporting errors or mismatches between the uploader and the person in the photo. We use the aligned version of the dataset and the official 5-fold split as recommended by [paplham2024call].
  • 2.
    Diabetic Retinopathy (DR) [diabetic-retinopathy-detection]: This medical imaging dataset contains over 88K high-resolution retinal fundus images for detecting and grading diabetic retinopathy. The task is to classify each image into one of 5 ordinal severity levels: (0: No DR, 1: Mild, 2: Moderate, 3: Severe, 4: Proliferative DR). The dataset exhibits significant challenges, including variations in imaging conditions and image quality. We use the official training/test split and further divide the training data into 5 folds for our cross-validation setup.
4.1.2 Evaluation Metrics
@@
4.1.3 Noise Injection Protocol
Label noise in ordinal tasks is typically asymmetric, as mislabelings are more likely to occur between adjacent, semantically similar classes. To simulate this realistically, we inject Gaussian Asymmetric Noise into the training and validation sets, following the protocol used in prior works [liu2024distributed, yuan2023cassor, jiang2024noise]. A noise transition matrix $T$, where $T_{ij}=P(\tilde{y}=j|y=i)$ is the probability of a true label $i$ being flipped to a noisy label $j$, is generated. The off-diagonal elements are defined by a Gaussian function of the distance between ranks:
@@ 4.1.3 Noise Injection Protocol
4.1.4 Implementation Details
All experiments use a ResNet-50 architecture, pre-trained on ImageNet, as the feature extractor backbone. For our proposed method, the model is trained for a maximum of 50 epochs ($E_{\text{max}}$), with the adaptive correction mechanism activating after 10 warm-up periods ($E_{\text{corr}}$). The base learning rates for correction were set to $\alpha_{\text{base}}=0.2$ and $\beta_{\text{base}}=0.8$ based on the hyperparameter analysis in 4.5.2.
4.2 Performance under Label Noise
We first evaluate the robustness of ORDAC against synthetic label noise. We compare our three proposed variants (ORDAC, ORDACC, and ORDACR) with two strong baselines: CORAL [cao2020rank], a standard ordinal regression method, and DLDL-v2 [gao2018age], a fixed-form LDL method. The training and validation sets are corrupted with Gaussian asymmetric noise at rates $\tau\in\{0.2,0.4\}$. The case $\tau=0$ corresponds to training on the original, uncorrupted datasets.
The results are presented in Table 1 and visualized in Figure 2. Our proposed methods consistently and significantly outperform the baselines across both datasets and all noise levels. As expected, the performance of all methods degrades as the noise rate increases. However, the ORDAC framework demonstrates substantially greater resilience. For instance, on Adience with 40% noise, ORDACR achieves an MAE of 0.6283, a dramatic improvement over CORAL (1.0427) and DLDL-v2 (0.8649).